# Selecting informative conformal prediction sets with false coverage rate control

### v1

@@ A.3 Examining Assumption 4
[…]
@@
Proposition A.2.
Let us consider a model where the variables $(X_{i},Y_{i})$, $i\in[n+m]$, are exchangeable and score functions satisfying Assumptions 6. Then the $p$-value family $\mathbf{\overline{p}}$ given by (7) satisfies Assumption 4.
@@
The proof is similar to the proof of Proposition A.2. We provide it below for completeness.
@@
A.4 Examining Assumption 5
Proposition A.4.
@@ A.4 Examining Assumption 5
Note that the above result is also true for selection rule of the type $\mathcal{S}(\mathbf{p})=\{i\in[m]\>:\>q_{i}\leq\tau\}$ for some fixed threshold $\tau$.
@@
Hence, the result is proved as soon as $\mathbf{q}$ is proved to be coordinate-wise nondecreasing in each $p$-value trajectory. This holds by Lemma A.5. ∎
Lemma A.5.
Suppose Assumption 2, then, almost surely, $\mathbf{q}$ defined by (11) is a nondecreasing function of each $p$-value ${p}^{(y)}_{i}$, $i\in[m]$ and $y\in\mathcal{Y}$.
@@
Let $\mathbf{p}$ and $\mathbf{p}^{\prime}$ be two $p$-value collections with ${p}^{(y)}_{i}\leq{p}^{\prime(y)}_{i}$, for all $i\in[m]$ and $y\in\mathcal{Y}$, with corresponding values $\mathbf{q}$ and $\mathbf{q}^{\prime}$. Let $i\in[m]$. By definition, $\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p}^{\prime})\in\mathcal{I}$ and $\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p})\subset\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p}^{\prime})$. By Assumption 2 (i) (iii), we have $\mathcal{C}^{q^{\prime}_{i}}_{n+i}(\mathbf{p})\in\mathcal{I}$, which in turn implies $q_{i}\leq q^{\prime}_{i}$ by definition of $q_{i}$. ∎
Appendix B Procedures using weighted class-calibrated $p$-values
Our procedures do not take into account the proportion of labels in each class in the test sample. However, in the classification case, these proportions are estimable from the data, and the estimates can aid inference.
Namely, let $\pi_{k}=\sum_{i=1}^{m}{\mathbf{1}\{Y_{n+i}=k\}}/m$ the true proportion of label $k$ in the test sample, $k\in[K]$, and consider the following possible estimates:
@@ Appendix B Procedures using weighted class-calibrated $p$-values
  • Storey-$\lambda$ estimator: $\hat{\pi}^{{\tiny\mbox{Storey}}}_{k}=\Big{(}1+\sum_{i=1}^{m}{\mathbf{1}\{p^{(k)}_{i}>\lambda\}}\Big{)}/(m(1-\lambda))$, $k\in[K]$.
The second one is similar to the classical estimator of true null hypotheses proportion in multiple testing (Storey,, 2002). Given the class-calibrated $p$-value family $\mathbf{{\widetilde{p}}}$ and one of the estimators $\hat{\pi}_{k}\in\{\hat{\pi}^{{\tiny\mbox{cal}}}_{k},\hat{\pi}^{{\tiny\mbox{Storey}}}_{k}\}$, we define the corresponding adaptive (weighted) $p$-value collection $\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt}}}=(\tilde{p}^{(k)}_{i,{\tiny\mbox{adapt}}},k\in[K],i\in[m]\}$ by
@@
Proposition B.1.
[…]
@@
Proposition B.1 is proved in § C.3. The bound (31) is only sharp when the labels are generated in the same way in the calibration and test sample (which implies the correct control in (ii)), so Adapt-InfoSP should not be used if the label proportions are expected to be (very) different between calibration and test samples.
We illustrate in Figure 9 the performance of the adaptive procedures for nonnull selection and non-trivial selection, respectively, in the set-up of unbalanced classes described in § 5.3. For the adaptive versions, $w_{k}=1/K$ for all $k\in[K]$, and $\lambda=1/2$. We consider two settings for the class conditional model: the test sample has class probabilities 0.1,0.1, 0.8 (i.e., a small label shift), and class probabilities 0.4,0.4,0.2 (i.e., a large label shift).
@@
Appendix C Proofs
C.1 Proof of Theorem 3.1
The proof is straightforward from the theory developed in § A. Namely, we apply Theorem A.1 by checking the two required assumptions: Assumption 4 holds for the two considered $p$-value collections (§ A.3); Assumption 5 holds for the considered $\mathrm{BH}(\mathbf{q})$ selection (§ A.4).
C.2 Proof of Theorem 3.3
The FCR control is a consequence of Theorem A.1: we only have to check that the pre-processed $p$-value family $\mathbf{\overline{p}}^{0}=(\overline{p}^{(0,y)}_{i},i+n\in\mathcal{S}^{(0)}\cap[n+1,n+m],y\in\mathcal{Y})$ (based on the conformal selection rule $\mathcal{S}^{(0)}$) satisfies Assumption 4 conditionally on $\mathcal{S}^{(0)}$. By Proposition A.2, it is sufficient to check that the variables $(X_{i},Y_{i})$, $i\in\mathcal{S}^{(0)}$, are exchangeable conditionally on $\mathcal{S}^{(0)}$. This is a direct consequence of the fact that the $(X_{j},Y_{j})$’s are iid and of the assumed permutation preserving property of $\mathcal{S}^{(0)}$.
@@ Appendix C Proofs
C.3 Proof of Proposition B.1
Let us first prove (i) by considering the class-conditional model. We follow the proof of Theorem A.1 (see § A.2) and we use that the $p$-value collection $\mathbf{{\widetilde{p}}}$ satisfies Assumption 4 (see Proposition A.3), where the probability in the super-uniform property (27) holds in the class-conditional model with ${\widetilde{W}}_{i}$, $i\in[m]$ given by (29). We also use that the informative selection rule $\mathcal{S}(\cdot)$ satisfies Assumption 5 (see Proposition A.4). Hence, following the same approach as in § A.2, and denoting $\mathbf{{\widetilde{p}}}_{{\tiny\mbox{cal}}}:=\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt-cal}}}$ and $\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}}:=(\tilde{p}^{(y)}_{j,{\tiny\mbox{cal}}})_{j\neq i,y\in\mathcal{Y}}$, we obtain
@@
Hence, $s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})$ can be written as a function $g(p^{(Y_{n+i})}_{i})$, with $g:u\mapsto s^{\min}_{i}(\Phi(Y,\Psi_{i}(u,W_{i})))$ nonincreasing and only depending on $Y$ and $W_{i}$. Applying Lemma D.1 for $c=w_{Y_{n+i}}/(m\hat{\pi}^{{\tiny\mbox{cal}}}_{Y_{n+i}})$, we obtain
@@
by using Lemma D.2 for the last inequality.
C.4 Proof of Proposition 5.1
@@
Now prove Proposition 5.1, by showing the equality for $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\breve{p}})$ and by carefully modifying the proof of Theorem A.1 (§ A.2) in order to get only equalities, by using that we consider the case of the non-trivial selection, that is, $\mathcal{S}=\mathrm{BH}(\mathbf{q})$, with $q_{i}=\min_{k\in[K]}\breve{p}^{(k)}_{i}$. By definition of the FCP (2), we have (remember also that the prediction set always includes $\arg\min_{k\in[K]}S_{k}(X_{n+i})$ due to the post-processing)
@@
(with equality if $(n+1)\alpha/m$ is an integer). By Lemma D.2, the expectation of the latter is equal to $\alpha(1-(1-\mathbb{P}(R_{1}>1))^{n+1})$.
@@
by letting $A_{i}=\{S_{Y_{k}}(X_{j}),j\in[n]:R_{k}>1\}\cup\{S_{Y_{n+i}}(X_{n+i})\}$ (all distinct almost surely). The third equality above is true because $K=2$ and $R_{n+i}>1$. We apply now Lemma D.3 because $\mathbf{q}=(q_{j},1\leq j\leq m)$ and $\mathbf{q}^{\prime}=(q^{\prime}_{j},1\leq j\leq m)$ (defined as above and with $q_{i}^{\prime}=1/(n+1)$) satisfy (36). Indeed, if $q_{j}>q_{i}$ for $j\neq i,$ then $S_{\max}(X_{n+j})\leq S_{\max}(X_{n+i})$ and $q_{j}=q^{\prime}_{j}$. Hence, we obtain, by letting $K_{i}=|\mathrm{BH}(\mathbf{q^{\prime}})|$ (note that $\mathbf{q^{\prime}}$ depends on $i$)

### v3

@@ C.3 Examining Assumption 4
For any $y\in\mathcal{Y}$; $S_{y}(\cdot)$ is of the form $S_{y}\big{(}\cdot\>;\>\mathcal{D}_{{\tiny\mbox{train}}},(X_{i})_{i\in[n+m]}\big{)}$ for an independent training data sample $\mathcal{D}_{{\tiny\mbox{train}}}$ and is invariant by permutation of the elements of $(X_{i})_{i\in[n+m]}$. In addition, the scores have no ties and the score function is regular in the sense of Assumption 1.
@@
Proposition C.2.
Let us consider a model where the variables $(X_{i},Y_{i})$, $i\in[n+m]$, are exchangeable conditionally on $\mathcal{D}_{{\tiny\mbox{train}}}$ and score functions satisfying Assumptions 6. Then the $p$-value family $\mathbf{\overline{p}}$ given by (7) satisfies Assumption 4.
@@
The proof is similar to the proof of Proposition C.2. We provide it below for completeness.
@@
C.4 Concordant selection rules
Proposition C.4.
@@ C.4 Concordant selection rules
Hence, the result is proved as soon as $\mathbf{q}$ is proved to be coordinate-wise nondecreasing in each $p$-value trajectory. This holds by Lemma C.5. ∎
Lemma C.5.
Suppose Assumption 2, then for each $i\in[m]$, almost surely, $q_{i}$ defined by (11) is a nondecreasing function of each ${p}^{(y)}_{i}$, $y\in\mathcal{Y}$.
@@
First note that each $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ only depends on the collection $\mathbf{p}=({p}^{(y)}_{i},y\in\mathcal{Y},j\in[m])$ through $({p}^{(y)}_{i},y\in\mathcal{Y})$, see (10), hence $q_{i}$ is only a function of $({p}^{(y)}_{i},y\in\mathcal{Y})$. Also, this allows to denote $\mathcal{C}^{\alpha}_{n+i}(\mathbf{p})$ simply by $\mathcal{C}^{\alpha}_{n+i}({p}^{(y)}_{i},y\in\mathcal{Y})$. Let $({p}^{(y)}_{i},y\in\mathcal{Y})$ and $({p}^{\prime(y)}_{i},y\in\mathcal{Y})$ be two $p$-value collections with ${p}^{(y)}_{i}\leq{p}^{\prime(y)}_{i}$, for all $y\in\mathcal{Y}$, with corresponding values $q_{i}$ and $q^{\prime}_{i}$. By definition, $\mathcal{C}^{q^{\prime}_{i}}_{n+i}({p}^{\prime(y)}_{i},y\in\mathcal{Y})\in\mathcal{I}$ and $\mathcal{C}^{q^{\prime}_{i}}_{n+i}({p}^{(y)}_{i},y\in\mathcal{Y})\subset\mathcal{C}^{q^{\prime}_{i}}_{n+i}({p}^{\prime(y)}_{i},y\in\mathcal{Y})$. By Assumption 2 (i) (iii), we have $\mathcal{C}^{q^{\prime}_{i}}_{n+i}({p}^{(y)}_{i},y\in\mathcal{Y})\in\mathcal{I}$, which in turn implies $q_{i}\leq q^{\prime}_{i}$ by definition of $q_{i}$.
Let us now discuss the concordance assumption for other, non-informative, selection rules. First note that Assumption 5 is true for any selection rule of the type $\mathcal{S}(\mathbf{p})=\{i\in[m]\>:\>q_{i}\leq t\}$ for some deterministic threshold $t$. More generally, let us consider the monotone $p$-value-based thresholding selection rule
$\mathcal{S}(\mathbf{p})=\{i\in[m]\>:\>f(p^{(y)}_{i},y\in\mathcal{Y})\leq t\},$
where $f:[0,1]^{\mathcal{Y}}\mapsto\mathbb{R}$ is some measurable coordinate-wise nondecreasing function. This selection rule satisfies Assumption 5. Indeed, we easily check in this case $\mathcal{A}(\mathbf{p}_{-i})=\{z\in[0,1]^{\mathcal{Y}}:f(z)\leq t\}$ (which does not depend on $\mathbf{p}_{-i}$) and thus
$s_{i}^{\min}(\mathbf{p}_{-i})=\Big{(}1+\sum_{j\neq i}{\mathbf{1}\{f(p^{(y)}_{j},y\in\mathcal{Y})\leq t\}}\Big{)}{\mathbf{1}\{\exists z\in[0,1]^{\mathcal{Y}}:f(z)\leq t\}},$
[…]
Finally, we mention that Assumption 5 is not satisfied in our context for top-$k$ type selection rules because of ties. Indeed, let us consider the simple case where $\mathcal{S}(\mathbf{p})$ selects the $2$-smallest $p$-values among the $p$-values $p_{i}:=p_{i}^{(y_{0})}$, $i\in[m]$, for some pre-specified $y_{0}\in\mathcal{Y}$ and $m=3$. When the $p$-values are at the minimum value $1/(n+1)$, the selection rule should make a unique choice and let us say that it selects $\{1,2\}$ when $(p_{1},p_{2},p_{3})=(1/(n+1),1/(n+1),1/(n+1))$. Then, for $i=3$, we easily obtain $s_{i}^{\min}(1/(n+1),1/(n+1),1/(n+1))=0$ (because $i=3$ is not selected), while $s_{i}^{\min}(1/(n+1),2/(n+1),1/(n+1))=2$ (because $i=3$ is selected). This contradicts the concordance assumption.
Nevertheless, by using a classical randomization trick for the conformal $p$-values Vovk et al., (2005); Bates et al., (2023); Jin and Candes, (2023) the $p$-value family has almost surely distinct values and top-$k$ type selection rules are concordant (under light additional assumptions), which concurs with the observation made in Benjamini and Bogomolov, (2013). More specifically, for $k\in[m]$, assume $\mathcal{S}(\mathbf{p})$ selects the $k$-smallest $f_{i}=f(p^{(y)}_{i},y\in\mathcal{Y})$ where $f:(0,1)^{\mathcal{Y}}\mapsto(0,1)$ is some measurable coordinate-wise nondecreasing surjective function and that the $f_{i}$’s have almost surely no ties. Then $\mathcal{A}(\mathbf{p}_{-i})=\{z\in[0,1]^{\mathcal{Y}}:i\in\mathcal{S}(z,\mathbf{p}_{-i})\}$ is not empty, because $z$ such that $f(z)<\min_{j\in[m]}f_{j}$ always belong to it. Hence, by definition $s_{i}^{\min}(\mathbf{p}_{-i})=k$ for all values of $\mathbf{p}_{-i}$, which establishes concordance. For instance, the concordance holds when using randomized conformal $p$-values and selecting the $k$-smallest $\mathcal{I}$-adjusted $p$-values, with $f(p^{(y)}_{i},y\in\mathcal{Y})=q_{i}$ being equal to $p_{i}^{(y_{0})}$ (excluding $y_{0}$), $\max_{y\in[a,b]}p_{i}^{(y)}$ (excluding $[a,b]$) or $\min_{k\in[K]}p_{i}^{(k)}$ (non-trivial classification).
Appendix D Procedures using weighted class-calibrated $p$-values
Procedure InfoSP does not take into account the proportion of labels in each class in the test sample. However, in the classification case, these proportions are estimable from the data, and the estimates can aid inference.
Let $\pi_{k}=\sum_{i=1}^{m}{\mathbf{1}\{Y_{n+i}=k\}}/m$ the true proportion of examples with label $k$ in the test sample, $k\in[K]$, and consider the following possible estimates:
@@ Appendix D Procedures using weighted class-calibrated $p$-values
  • Storey-$\lambda$ estimator: $\hat{\pi}^{{\tiny\mbox{Storey}}}_{k}=\Big{(}1+\sum_{i=1}^{m}{\mathbf{1}\{p^{(k)}_{i}>\lambda\}}\Big{)}/(m(1-\lambda))$, $k\in[K]$. It is similar to the classical estimator of true null hypotheses proportion in multiple testing (Storey,, 2002).
Given the class-calibrated $p$-value family $\mathbf{{\widetilde{p}}}$ and one of the estimators $\hat{\pi}_{k}\in\{\hat{\pi}^{{\tiny\mbox{cal}}}_{k},\hat{\pi}^{{\tiny\mbox{Storey}}}_{k}\}$, we define the corresponding adaptive (weighted) $p$-value collection $\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt}}}=(\tilde{p}^{(k)}_{i,{\tiny\mbox{adapt}}},k\in[K],i\in[m]\}$ by
@@
Proposition D.1.
Consider an informative subset collection $\mathcal{I}$ satisfying Assumption 2, score functions satisfying Assumption 1 and the $p$-value collection $\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt-cal}}}=(\tilde{p}^{(k)}_{i,{\tiny\mbox{adapt-cal}}},k\in[K],i\in[m]\}$ defined as in (28) with the calibration-based estimator $\hat{\pi}^{{\tiny\mbox{cal}}}_{k}=(|\mathcal{D}^{(k)}_{{\tiny\mbox{cal}}}|+1)/(n+1),$ $k\in[K]$. Then the corresponding Adapt-InfoSP procedure $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt-cal}}})$ satisfies the following:
@@
Proposition D.1 is proved in § E.3. The bound (29) is only sharp when the labels are generated in the same way in the calibration and test sample (which implies the correct control in (ii)), so Adapt-InfoSP should not be used if the label proportions are expected to be (very) different between calibration and test samples.
We illustrate in Figure 6 the performance of the adaptive procedures for nonnull selection and non-trivial selection, respectively, in the set-up of unbalanced classes described in § 5.2. For the adaptive versions, $w_{k}=1/K$ for all $k\in[K]$, and $\lambda=1/2$. We consider two settings for the class-conditional model: the test sample has class probabilities 0.1,0.1, 0.8 (i.e., a small label shift), and class probabilities 0.4,0.4,0.2 (i.e., a large label shift).
@@
In the simulations we carried out for the iid settings, Adapt-InfoSP on class-conditional $p$-values had worse power than InfoSCOP (omitted for brevity).
@@
Appendix E Proofs
E.1 Proof of Theorem 3.1
The proof is straightforward from the theory developed in § C. Namely, we apply Theorem C.1 by checking the two required assumptions: Assumption 4 holds for the two considered $p$-value collections (§ C.3); Assumption 5 holds for the considered $\mathrm{BH}(\mathbf{q})$ selection (§ C.4). The result thus follows from the fact that by Proposition C.4 we have
$\mathcal{R}_{\alpha}=\big{(}\mathcal{C}^{\alpha(1\vee s_{i}^{\min})/m}_{n+i}(\mathbf{p})\big{)}_{i\in\mathcal{S}}=\big{(}\mathcal{C}^{\alpha|\mathrm{BH}(\mathbf{q})|/m}_{n+i}(\mathbf{p})\big{)}_{i\in\mathrm{BH}(\mathbf{q})}=\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{p}).$
E.2 Proof of Theorem 3.3
The proof is a consequence of Lemma H.5 applied with the FCR criterion: condition (i) in Lemma H.5 is satisfied from Theorem 3.1 (which is true more broadly in the case of exchangeable samples, see Theorem C.1 and Proposition C.2); condition (ii) in Lemma H.5 follows from the assumed permutation preserving property of $\mathcal{S}^{(0)}$. Hence, the conclusion of Lemma H.5 applies which gives the FCR control of InfoSCOP.
@@ Appendix E Proofs
E.3 Proof of Proposition D.1
Let us first prove (i) by considering the class-conditional model. We follow the proof of Theorem C.1 (see § C.2) and we use that the $p$-value collection $\mathbf{{\widetilde{p}}}$ satisfies Assumption 4 (see Proposition C.3), where the probability in the super-uniform property (25) holds in the class-conditional model with ${\widetilde{W}}_{i}$, $i\in[m]$ given by (27). We also use that the informative selection rule $\mathcal{S}(\cdot)$ satisfies Assumption 5 (see Proposition C.4). Hence, following the same approach as in § C.2, and denoting $\mathbf{{\widetilde{p}}}_{{\tiny\mbox{cal}}}:=\mathbf{{\widetilde{p}}}_{{\tiny\mbox{adapt-cal}}}$ and $\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}}:=(\tilde{p}^{(y)}_{j,{\tiny\mbox{cal}}})_{j\neq i,y\in\mathcal{Y}}$, we obtain
@@
Hence, $s^{\min}_{i}(\mathbf{{\widetilde{p}}}_{-i,{\tiny\mbox{cal}}})$ can be written as a function $g(p^{(Y_{n+i})}_{i})$, with $g:u\mapsto s^{\min}_{i}(\Phi(Y,\Psi_{i}(u,W_{i})))$ nonincreasing and only depending on $Y$ and $W_{i}$. Applying Lemma H.1 for $c=w_{Y_{n+i}}/(m\hat{\pi}^{{\tiny\mbox{cal}}}_{Y_{n+i}})$, we obtain
@@
by using Lemma H.3 for the last inequality.
E.4 Proof of Proposition 5.1
@@
Now prove Proposition 5.1, by showing the equality for $\mathcal{R}_{\alpha}^{{\tiny\textnormal{InfoSP}}}(\mathbf{\breve{p}})$ and by carefully modifying the proof of Theorem C.1 (§ C.2) in order to get only equalities, by using that we consider the case of the non-trivial selection, that is, $\mathcal{S}=\mathrm{BH}(\mathbf{q})$, with $q_{i}=\min_{k\in[K]}\breve{p}^{(k)}_{i}$. By definition of the FCP (2), we have (remember also that the prediction set always includes $\arg\min_{k\in[K]}S_{k}(X_{n+i})$ due to the post-processing)
@@
(with equality if $(n+1)\alpha/m$ is an integer). By Lemma H.3, the expectation of the latter is equal to $\alpha(1-(1-\mathbb{P}(R_{1}>1))^{n+1})$.
@@
by letting $A_{i}=\{S_{Y_{k}}(X_{j}),j\in[n]:R_{k}>1\}\cup\{S_{Y_{n+i}}(X_{n+i})\}$ (all distinct almost surely). The third equality above is true because $K=2$ and $R_{n+i}>1$ and thus $S_{Y_{n+i}}(X_{n+i})=S_{\max}(X_{n+i})$. We apply now Lemma H.4 because $\mathbf{q}=(q_{j},1\leq j\leq m)$ and $\mathbf{q}^{\prime}=(q^{\prime}_{j},1\leq j\leq m)$ (defined as above and with $q_{i}^{\prime}=1/(n+1)$) satisfy (34). Indeed, if $q_{j}>q_{i}$ for $j\neq i,$ then $S_{\max}(X_{n+j})\leq S_{\max}(X_{n+i})$ and $q_{j}=q^{\prime}_{j}$. Hence, we obtain, by letting $K_{i}=|\mathrm{BH}(\mathbf{q^{\prime}})|$ (note that $\mathbf{q^{\prime}}$ depends on $i$)